A sphere of radius 1.00 cm is placed in the path of a parallel beam of light of large aperture. The intensity of the light is 0.50 W/cm^2. If the sphere completely absorbs the radiation falling on it, find the force exerted by the light beam on the sphere.

Grade:upto college level

1 Answers

Kevin Nash
9 years ago
Sol. We know, If a perfectly reflecting solid sphere of radius ‘r’ is kept in the path of a parallel beam of light of large aperture if intensity is I, force = πr^2l/C l = 0.5 W/c^2, -= 1 cm, C = 3 * 10^8 m/s force = π *(1)^2 *0.5/3 * 10^8 3.14 * 0.5/3 * 10^8/3 * 10^8 = 0.523 * 10^-8 = 5.2 * 10^-19 N.